# How Haag-tied is QFT, really?

### 18 Dec 2023 v2

@@ 4.5 Maiezza and Vasquez
In Maiezza and Vasquez’s assessment, Haag’s theorem reveals a central problem when we try to improve upon the standard methods of perturbatively renormalized canonical QFT. We see this problem clearly when we consider the entire perturbative expansion series. The problem stems from vacuum polarization: because of vacuum polarization, the full power-series expansion diverges [^34] This was first noted in [Dyson, 1952] concerning QED.. Maiezza and Vasquez prove that renormalon singularities arise in the total perturbative series, and trace these back to Haag’s theorem. These renormalon singularities “are the concrete manifestation” of the impossibility of generating an unambiguous finite result in such cases [Maiezza and Vasquez, 2020, page 10].  [^35] Renormalons are singularities that arise as a function of the complex transform parameter when a formally divergent series is summed using Borel summation, see [Beneke, 1999] for a review. Whilst Haag’s theorem is generally associated with infrared divergences of individual terms in the perturbative series (see [Earman and Fraser, 2006, p. 319], [Fraser, 2006, pp 63-66], [Sbisà, 2020, p. 34], [Van Hove, 1952] ), Maiezza and Vasquez associate the overall divergence of the perturbative series back to Haag’s theorem. As such, Maiezza and Vasquez associate both IR (low momentum) and UV (high momentum) renormalon divergences with Haag’s theorem.
@@
According to Maiezza and Vasquez, there is no repair we can make at this time to address this problem. There remains a concerning flaw in perturbative renormalization, namely, the renormalized series’ dependence on an arbitrary choice of constant. As Maiezza and Vasquez put it, perturbative renormalization “cannot be a self-consistent cure, because perturbative renormalization needs to be completed, or in practice resummed” [Maiezza and Vasquez, 2020, page 10].
@@
That said, Maiezza and Vasquez are quick to note that the practical uses made of the interaction picture with perturbative renormalization need no repair: they evade Haag’s theorem. Haag’s theorem is, after all, a non-perturbative result, in that it applies to the entire power-series expansion, not just the perturbative expansion at some finite order. Because the power-series expansion is asymptotic, truncating at finite orders can give meaningful approximations, while also missing out on the non-perturbative effects that lead to the complete series’ divergence. Thus, while regularization and renormalization do not ‘fix’ the problem posed by the divergences tracked by Haag’s theorem, this is essentially because Haag’s theorem does not apply in cases where the series is truncated. Perturbative renormalization fixes the divergences of individual terms within the series, but it does not address the non-perturbative divergences of Haag’s theorem.
@@
They suggest two forms of renovation. First, in the short run, research efforts should aim at improving our understanding of the non-perturbative singularities that arise, such as renormalons. Further research should be directed into mathematical techniques to better understand neglected non-perturbative effects (for example [Maiezza and Vasquez, 2021b, Maiezza and Vasquez, 2021a]). These include resurgence, a method for reading off certain non-perturbative effects from the perturbative expansions (for a review, see [Dunne and Ünsal, 2016]).
@@ 4.6 Kastner
[Kastner, 2015]’s outlook differs from the previous authors. Insofar as she seeks the correct formulation of our theory of quantum phenomena, she sits with the physicists above. However, insofar as she thinks this correct formulation requires a fundamental reinterpretation of quantum phenomena, she sits with the theory interpreters, who are overwhelmingly philosophers. Kastner advocates replacing QFT with a direct action theory (DAT). As such, she has already given up on mediating interactions locally in favor of nonlocal, direct interactions between field sources. This is a profound departure from QFT.
@@
Kastner locates the problem ‘further back’ conceptually. Rather, Kastner begins with the commitment that QFT is wrong from the outset. As such, her assessment of QFT in the light of Haag’s theorem is grim: “QFT is not the correct model; a different, yet empirically equivalent, model is needed” (57). The problem lies with the foundational assumption of QFT—namely, that we use fields as mediators between particles (crudely, as “a ‘bucket brigade’ that is invoked in order to restrict causal influence to a local, continuous conveyance from spacetime point to spacetime point” (59)). Thus, Haag’s theorem “simply tells us what we already know: the interaction picture of quantized fields does not really exist” [Kastner, 2015, 58]. Haag’s theorem is therefore an additional motivation for abandoning QFT in favor of DATs.
@@
The treatment for such a deep problem is a major change in the modeling procedure for this area of physics. If we accept that Haag’s theorem shows us that it simply does not work to model interactions by field operators creating and destroying Fock space states, then, Kastner urges, we are to instead revive direct action theories (DAT), in which we model interactions through a direct, nonlocal interaction between sources of the field [Kastner, 2015, 57].
Rather than modeling the evolution from free to interacting states (as in the interaction picture of QFT), on the direct action approach, we do away with fields as mediators of interactions. Instead, we posit a direct, nonlocal between particles or sources of the field, without needing any intermediary field, so that the issue of constructing a unitary operator to map between free and interacting theories does not arise. More fundamentally, in a DAT, quantum mechanics is only incorporated at the level of the sources and receivers of a field, and is not enforced through the imposition of canonical commutation relations on the fields (see [Davies, 1970]). As such, there is no requirement of local commutativity between the field operators at spacelike separations, unlike in ordinary QFT (recall axiom 3).
Kastner believes that this is a viable repair because she is convinced that DATs—which differ radically from QFTs—are empirically equivalent to QFTs. This places significant weight on (purported) demonstrations of the equivalence of direct-action theories and QFT (e.g., [Narlikar, 1968]). If the demonstrations are sound, and they extend such that all currently-used QFTs are approximations to some DAT, then they would go some way towards explaining how calculations in renormalized quantum field theories are able to generate successful predictions: QFT’s “mathematical inconsistencies can be rendered inconsequential since they can be understood as arising from its ‘makeshift,’ nonfundamental character” [63]. [^37] However, this is a non-trivial ‘if’, especially given the recent results suggesting that ‘empirical equivalence’ is a less straightforward concept than philosophers often presume. See [Weatherall, 2019] and references therein for more detail.
@@ 4.7 Seidewitz
Seidewitz’s outlook is interesting in that it substantially coincides with several of the others already discussed, while at the same time departing from them. First, he seems to think of theory interpretation in a way similar to Earman and D. Fraser and the theory interpreters. Like the latter, he appears to take mathematical rigor as a precondition for providing a satisfactory interpretation [Seidewitz, 2017, 369]. However, Seidewitz sees axiomatic QFT built on the Wightman axioms as an incorrect marriage of special relativity with quantum theory. In the light of this, Seidewitz thinks that these axioms will need to face significant revision.  [^39] See [Fraser, 2006, Section 5] for an extended discussion of the sense in which Fraser takes the axioms of QFT to be provisional. Seidewitz departs from the (pristine) theory interpreters insofar as a univocal characterization of particle-ness is not forthcoming. This is because particles now come in two varieties—virtual and real—and a new definition of ‘particle’ will have to account for the significant mathematical distinction between the two. [^40] In particular, we should expect (an analog of) the (global) number operator characterization to hold only for real particles, given that the space-like evolution possible for virtual particles prevents their characterization as Fock space states to be created or destroyed (where this is done by operators used to define number operators).
@@
As Seidewitz sees it, Haag’s theorem “essentially relies on a conflict between the presumption that the fields are Lorentz-covariant and the special identification of time in the assumptions of Theorem 1” [Seidewitz, 2017, 360]. This special identification is apparent in that the time-evolution is given by the (frame-dependent) Hamiltonian. This requires particles to evolve on time-like trajectories. Essentially, the (frame-dependent) Hamiltonian operator is playing double-duty as the generator of time translation, in addition to the generator of state evolution, with $t$ as the evolution parameter for each. However, because of this, the translation group of $t$—the unitary operators $\hat{G}(t)$—coincides with the group of Lorentz transformations, which guarantees that one can extend the coincidence $\hat{\psi}_{2}(t,\mathbf{x})=\hat{G}\hat{\psi}_{1}(t,\mathbf{x})\hat{G^{-1}}$ at a specific time $t$ to all times $t$.
@@
Then, one can consider the one-dimensional group generated by each Hamiltonian, where $\lambda$ (instead of $t$) is the (now frame-independent!) evolution parameter. (Physically, $\lambda$ can be thought of as the parameterization of a particle’s path in spacetime—now unburdened from the constraint that the path be timelike.) Haag’s theorem is evaded by limiting the equality of vacuum expectation values (Theorem 1) to equal values of $\lambda$, which is neither surprising—each Hamiltonian has a unique Poincaré invariant vacuum state—nor problematic—one can no longer use Lorentz transformations to extend the equality any farther. In particular, one cannot demonstrate that the fields coincide (Haag’s theorem). Thus, interacting fields need not coincide with free fields, and yet a version of the interaction picture is reinstated since interacting fields can be related to free fields by a unitary intertwiner [Seidewitz, 2017, 369].
@@ 4.8 Framework Results Summary
The viewpoints above are characterized by several different purposes. Duncan and Miller are both concerned with a specific explanatory task: how, in fact, do physical predictions evade Haag’s theorem? This takes for granted the acceptability of current predictions and appropriateness of QFT’s scale-relative form. While Klaczynski makes similar observations, he moreover desires a QFT form making the algebraic structure of interaction dynamics mathematically manifest. Similarly, Maiezza and Vasquez hope to precisely characterize the non-perturbative structure of QFT. Klaczynski and Maiezza and Vasquez therefore coincide with Duncan and Miller in their attention to canonical QFT, but diverge in setting their sights on an unambiguous and non-perturbative structure underlying it. In contrast, Kastner and Seidewitz divert their attention from canonical QFT altogether in search of a more satisfactory marriage of special relativity and quantum theory. This search beyond canonical QFT is shared by many theory interpreters, who seek a QFT form that straightforwardly exposes its ontology. Among these, the pristine interpreters’ search is allied with Earman and D. Fraser, insofar as they expect a straightforward exposure of QFT’s ontology will proceed from a theory whose consequences are logically-straightforward derivations from conceptually clear assumptions.
@@
Pristine Interpreters
Should we interpret QFT interactions as the interaction of (free-field) particles? Only the formal variants of QFT are ripe for interpretation. The process by which predictions are actually generated is irrelevant.
Non-Pristine Interpreters
@@
How should QM be extended to the phenomena described by QFT? We desire a well-defined theory eschewing the CCRs whose predictions coincide with those of canonical QFT.
@@ 5 Where and how to make future progress
Our ultimate goal has been to provide guidance on QFT’s construction in light of Haag’s theorem. Section 3 suggested that this would not be easy, given the literature’s apparent labyrinthine set of disagreements. Deploying our framework in 4 regimented and catalogued the various responses to Haag’s theorem in such a way that those actively working on the philosophy, mathematics, or physics of QFT can more readily appreciate where the real differences between these proposals lie. We now turn to giving our ‘general contractor’s’ viewpoint on the insight Haag’s theorem gives us into QFT, present and future. Section 5.1 uses the results of the framework application to disentangle three separate lines of debate on Haag’s theorem. In each of these debates, there is a clear guiding question to the discussion, such that debates are not entirely talking past each other. Nevertheless, we highlight for each debate the role that extra-Haagian outlooks continue to play in shaping the discourse. Our contribution is to sort out and elucidate the driving forces in the various debates around Haag’s theorem, without ourselves defending a position in any of those debates. However, in section 5.2, we take a clear stand on the general methodological lessons to be learned from this study of Haag’s theorem. We discuss prospects for future progress on the challenges for QFT raised by Haag’s theorem. We argue that stakeholders in physics, mathematics, and philosophy ought to give more specific characterizations of their proposed methods for making progress, together with more detailed arguments as to why those preferred methods are likely to success.

### 06 Apr 2024 v3

@@ 4.5 Maiezza and Vasquez
In Maiezza and Vasquez’s assessment, Haag’s theorem reveals a central problem when we try to improve upon the standard methods of perturbatively renormalized canonical QFT. Haag’s theorem is generally associated with infrared divergences of individual terms in the perturbative series (see [Earman and Fraser, 2006, p. 319], [Fraser, 2006, pp 63-66], [Sbisà, 2020, p. 34] ). However, Maiezza and Vasquez trace the overall divergence of the perturbative series back to Haag’s theorem. After all, “Haag’s theorem is non-perturbative and thus, for a finite value of the coupling, perturbative renormalization cannot be its solution” [Maiezza and Vasquez, 2020, page 7].
We see this problem clearly when we consider the entire perturbative expansion series. The problem stems from vacuum polarization: because of vacuum polarization, the full power series expansion diverges [^32] This was first noted in [Dyson, 1952] concerning QED.. Maiezza and Vasquez demonstrate that renormalon singularities arise in the total perturbative series, and trace these back to Haag’s theorem. These renormalon singularities “are the concrete manifestation” of the impossibility of generating an unambiguous finite result in such cases [Maiezza and Vasquez, 2020, page 10]. 333333 Renormalons are singularities that arise as a function of the complex transform parameter when a formally divergent series is summed using Borel summation, see [Beneke, 1999] for a review. Infrared renormalons are generally associated with long-distance or low energy non-perturbative effects, emerging from the accumulation of low-momentum virtual particles in loop corrections, whilst ultraviolet renormalons are generally associated with artefacts introduced by the renormalization procedure itself (see also [Parisi, 1979]). For non-asymptotically free models, Maiezza and Vasquez associate the UV renormalon divergences with Haag’s theorem. They argue that the same would hold for IR renormalons in asymptotically free models. [^34] Roughly speaking, in an asymptotically free quantum field theory, the interaction strength becomes weaker as the distance scale decreases or the energy scale increases. See [Politzer, 1973, Gross and Wilczek, 1973, David Politzer, 1974, Wilczek, 2005] for further details. Thus, either kind of renormalon divergence can be traced back to Haag’s theorem in some cases.  [^35] Maiezza and Vasquez’ make no de facto distinction between UV and IR renormalons (c.f. [Parisi, 1979]), instead showing how renormalons emerge from the renormalization group equations within the resurgence framework, without reference to specific Feynman diagrams (see also [Maiezza and Vasquez, 2021b, Maiezza and Vasquez, 2021a, Dunne and Ünsal, 2016]). As such, either UV or IR renormalons may be interpreted as a symptoms of Haag’s theorem, depending on the theory. For instance, in an (asymptotically free) Yang-Mills theory, IR renormalons would correspond to singularities in the semi-positive axis of the Borel plane, and render a Borel-Laplace resummation ambiguous. In this case, Maiezza and Vasquez argue that such IR renormalons can be traced back to Haag’s theorem.
@@
According to Maiezza and Vasquez, there is no straightforward way to repair canonical QFT to address this problem. There remains a concerning flaw in perturbative renormalization, namely, the renormalized series’ dependence on an arbitrary choice of constant. As Maiezza and Vasquez put it, perturbative renormalization “cannot be a self-consistent cure, because perturbative renormalization needs to be completed, or in practice resummed” [Maiezza and Vasquez, 2020, page 10].
@@
With that said, Maiezza and Vasquez are quick to note that the practical uses made of the interaction picture with perturbative renormalization need no repair: they evade Haag’s theorem. Haag’s theorem is, after all, a non-perturbative result, in that it applies to the entire power series expansion, not just the perturbative expansion at some finite order. Because the power series expansion is asymptotic, truncating at finite orders can give meaningful approximations, while also missing out on the non-perturbative effects that lead to the complete series’ divergence. Thus, while regularization and renormalization do not ‘fix’ the problem posed by the divergences tracked by Haag’s theorem, this is essentially because Haag’s theorem does not apply in cases where the series is truncated. Maiezza and Vasquez hold that perturbative renormalization fixes the divergences of individual terms within the series, but it does not address the non-perturbative divergences of Haag’s theorem. 373737[Miller, 2018] and Maiezza and Vasquez agree about how perturbative renormalization works to fix the divergences of individual terms; however, Miller does not connect this to the overall divergence of the perturbative series).
@@
They suggest three forms of renovation. First, in the short run, research efforts should aim at improving our understanding of the non-perturbative singularities that arise, such as renormalons. Further research should be directed into mathematical techniques to better understand neglected non-perturbative effects (for example [Maiezza and Vasquez, 2021b, Maiezza and Vasquez, 2021a]). These include resurgence, a method for reading off certain non-perturbative effects from the perturbative expansions (for a review, see [Dunne and Ünsal, 2016]).
@@
Third, Maiezza and Vasquez suggest physicists might need to consider more radical models, perhaps rejecting some of the standard assumptions of canonical QFT. They suggest one such possible model, dimensionally-reduced QFT (DRQFT) [Maiezza and Vasquez, 2023], in which the spacetime dimensions are dependent on the energy scale. Such an approach requires the introduction of a new mass-energy scale to encode some further unknown physics at high energy scales.  [^38] As such, this is not intended as a candidate for something like a “fundamental theory”. Motivated by approaches in quantum gravity, they interpret the scale as encoding the effects of a classical field of geometric origins that weights long and short contributions differently. So the sense in which unknown physics is encoded in this model is quite different from the approach in Wilsonian renormalization. However, something analogous to the interaction picture can be salvaged in such a theory. They demonstrate that such a theory can be made finite in perturbation theory, free of UV renormalon ambiguities. The key is that such a theory avoids one of the assumptions of Haag’s theorem: the vacuum states of DRQFT are not Poincaré-invariant (see [Maiezza and Vasquez, 2023, pages 17-18]).
@@ 4.6 Kastner
[Kastner, 2015]’s outlook differs from the previous authors. Insofar as she seeks the correct formulation of our theory of quantum phenomena, she sits with the physicists above. However, she thinks this correct formulation requires a fundamental reinterpretation of quantum phenomena. Here she goes further than the theory interpreters, who focus on new interpretations of existing physics. Kastner advocates replacing QFT with a direct-action theory (DAT). As such, she has already given up on mediating interactions locally in favor of nonlocal, direct interactions between field sources. This is a profound departure from QFT.
@@
Kastner locates the problem further back conceptually. Rather, Kastner begins with the commitment that QFT is wrong from the outset. As such, her assessment of QFT in the light of Haag’s theorem is grim: “QFT is not the correct model; a different, yet empirically equivalent, model is needed” (57). The problem lies with the foundational assumption of QFT—namely, that we use fields as mediators between particles (crudely, as “a ‘bucket brigade’ that is invoked in order to restrict causal influence to a local, continuous conveyance from spacetime point to spacetime point” (59)). Thus, Haag’s theorem “simply tells us what we already know: the interaction picture of quantized fields does not really exist” [Kastner, 2015, 58]. Haag’s theorem is therefore an additional motivation for abandoning QFT in favor of DATs.
@@
The treatment for such a deep problem is a major change in the modeling procedure for this area of physics. If we accept that Haag’s theorem shows us that it simply does not work to model interactions by field operators creating and destroying Fock space states, then, Kastner urges, we are to instead revive direct-action theories (DAT), in which we model interactions through a direct, nonlocal interaction between sources of the field [Kastner, 2015, 57].
Rather than modeling the evolution from free to interacting states (as in the interaction picture of QFT), on the direct-action approach, we do away with fields as mediators of interactions. Instead, we posit a direct, nonlocal between particles or sources of the field, without needing any intermediary field, so that the issue of constructing a unitary operator to map between free and interacting theories does not arise. More fundamentally, in a DAT, quantum mechanics is only incorporated at the level of the sources and receivers of a field, and is not enforced through the imposition of canonical commutation relations on the fields (see [Davies, 1970]). As such, there is no requirement of local commutativity between the field operators at spacelike separations, unlike in ordinary QFT (recall axiom 3).
Kastner believes that this is a viable repair because she is convinced that DATs—which differ radically from QFTs—are empirically equivalent to QFTs. This places significant weight on (purported) demonstrations of the equivalence of direct-action theories and QFT (e.g., [Narlikar, 1968]). If the demonstrations are sound, and they extend such that all currently used QFTs are approximations to some DAT, then they would go some way towards explaining how calculations in renormalized quantum field theories are able to generate successful predictions: QFT’s “mathematical inconsistencies can be rendered inconsequential since they can be understood as arising from its “makeshift,’ nonfundamental character” [63]. [^39] However, this is a non-trivial if, especially given the recent results suggesting that empirical equivalence is a less straightforward concept than philosophers often presume. See [Weatherall, 2019] and references therein for more detail.
@@ 4.7 Seidewitz
Seidewitz’s outlook is interesting in that it substantially coincides with several of the others already discussed, while at the same time departing from them. First, he seems to think of theory interpretation in a way similar to Earman and D. Fraser and the theory interpreters. Like the latter, he appears to take mathematical rigor as a precondition for providing a satisfactory interpretation [Seidewitz, 2017, 369]. However, Seidewitz sees axiomatic QFT built on the Wightman axioms as an incorrect marriage of special relativity with quantum theory. In the light of this, Seidewitz thinks that these axioms will need to face significant revision.  [^41] See [Fraser, 2006, Section 5] for an extended discussion of the sense in which Fraser takes the axioms of QFT to be provisional. Seidewitz departs from the puzzle-focused interpreters insofar as a univocal characterization of particle-ness is not forthcoming. This is because particles now come in two varieties—virtual and real—and a new definition of ‘particle’ will have to account for the significant mathematical distinction between the two. [^42] In particular, we should expect (an analog of) the (global) number operator characterization to hold only for real particles, given that the space-like evolution possible for virtual particles prevents their characterization as Fock space states to be created or destroyed (where this is done by operators used to define number operators).
@@
As Seidewitz sees it, Haag’s theorem “essentially relies on a conflict between the presumption that the fields are Lorentz-covariant and the special identification of time in the assumptions of Theorem 1” [Seidewitz, 2017, 360]. This special identification is apparent in that the time-evolution is given by the (frame-dependent) Hamiltonian. This requires particles to evolve on time-like trajectories. Essentially, the (frame-dependent) Hamiltonian operator is playing double duty as the generator of time translation, in addition to the generator of state evolution, with $t$ as the evolution parameter for each. However, because of this, the translation group of $t$—the unitary operators $\hat{G}(t)$—coincides with the group of Lorentz transformations, which guarantees that one can extend the coincidence $\hat{\psi}_{2}(t,\mathbf{x})=\hat{G}\hat{\psi}_{1}(t,\mathbf{x})\hat{G^{-1}}$ at a specific time $t$ to all times $t$.
@@
Then, one can consider the one-dimensional group generated by each Hamiltonian, where $\lambda$ (instead of $t$) is the (now frame-independent!) evolution parameter. (Physically, $\lambda$ can be thought of as the parameterization of a particle’s path in spacetime—now unburdened from the constraint that the path be timelike.) Haag’s theorem is evaded by limiting the equality of vacuum expectation values (Theorem 1) to equal values of $\lambda$, which is neither surprising—each Hamiltonian has a unique Poincaré-invariant vacuum state—nor problematic—one can no longer use Lorentz transformations to extend the equality any farther. In particular, one cannot demonstrate that the fields coincide (Haag’s theorem). Thus, interacting fields need not coincide with free fields, and yet a version of the interaction picture is reinstated since interacting fields can be related to free fields by a unitary intertwiner [Seidewitz, 2017, 369].
@@ 4.8 Framework Results Summary
The viewpoints above are characterized by several different purposes. We schematically organize these purposes in figure 1. Duncan and Miller are both concerned with a specific explanatory task: how, in fact, do physical predictions evade Haag’s theorem? This takes for granted the acceptability of current predictions and appropriateness of QFT’s scale-relative form. While Klaczynski makes similar observations, he moreover desires a QFT form that makes the algebraic structure of interaction dynamics mathematically manifest. Similarly, Maiezza and Vasquez hope to precisely characterize the non-perturbative structure of QFT. Klaczynski and Maiezza and Vasquez therefore coincide with Duncan and Miller in their attention to canonical QFT, but diverge in setting their sights on an unambiguous and non-perturbative structure underlying it. In contrast, Kastner and Seidewitz divert their attention from canonical QFT altogether in search of a more satisfactory marriage of special relativity and quantum theory.Finally, Earman and Fraser and the theory interpreters are doing the distinctly philosophical work of sorting out the downstream interpretive implications of Haag’s theorem in conversation with the wider goals and conceptual foundations of QFT.
@@
Theory Interpreters: constrained puzzle
Should we interpret QFT interactions as the interaction of (free field) particles? Only the formal variants of QFT are ripe for interpretation. The calculational details by which predictions are actually generated is irrelevant.
Theory Interpreters: beyond the puzzle
@@
How should QM be extended to the phenomena described by QFT? We desire a well-defined theory whose predictions coincide with those of canonical QFT, but which eschews the CCRs.
@@ 5 Where and how to make future progress
Our ultimate goal has been to provide guidance on QFT’s construction in light of Haag’s theorem. So what guidance does the framework provide? The reader might expect us to promote a QFT pluralism along the following lines. Section 3 suggested there were substantial disagreements about how, or even whether, the QFT apparatus works. These disagreements appeared to concern the propriety of our present theory. Had the scrutiny of setion 4 born this out, adjudication would have appeared necessary to ensure the security of our present theory. But this was not the upshot of section 4. Instead, our framework revealed that respondents’ purposes were not uniform. Moreover, the response-pluralism was decidedly future-directed in that the divergences were easier to appreciate when viewing the proposed renovations. In this way, different beliefs about how to pursue a future theory may be driving the differing assessment’s of Haag’s theorem. We do not even know what this future theory is, let alone how we will discover it. Thus, humility would suggest leaving each of these researchers to their own pursuits.
We find this pluralistic response tempting, and we think there is much right in it. Nevertheless, we do not think it is sufficient for providing guidance on QFT’s construction. In our opening vignette, we imagined researchers encountering Haag’s theorem, and we asked how they should respond. While a reader could have, understandably, chalked it up as mere rhetoric, we take the vignette seriously: how should actual researchers respond to the theorem? Actual researchers and actual research communities are subject to resource constraints, and they are thus forced to make choices. Indeed, some of these choices are difficult, in the sense that there are few and fallible reasons to prefer one choice over another. Yet such choices are made, and reasons for these difficult choices are given. Consider that each respondent we survey not only pursues their own path, but moreover defends their decision to do so. While pluralism is right in that there appear to be multiple reasonable choices, it does not help us human decision-makers reason through the choices we actually face.
But we also do not think methodological monism—picking just a single path—is necessary for guidance at this stage. That is, we do not think the framework must force a choice upon researchers in order to provide guidance to them. Rather, we take it that any information that helps researchers make the difficult decision(s) of what path(s) to pursue constitutes guidance. Here we content ourselves with two kinds of guidance. In 5.1, we use the framework to separate the wheat of 3 from the chaff by highlighting the very real choices of goal that underlie the more superficial disagreements. The above respondents have and provide reasons for these choices, and, given that the future of QFT is uncertain, these reasons are defeasible. Researchers are thus guided by the framework to choices of consequence to the future of physics and, ipso facto, to reasoning of consequence.
In 5.2, we show how the framework, more generally, guides researchers from more traditional philosophical questions about our foundations to more methodological ones. Traditionally, philosophers have asked pointed questions about the current status of specific concepts or claims—questions about rigor and mathematical existence, about theory interpretation, about theory essence. But as section 5.1 shows, these traditional questions morph into questions about what kinds of reasoning we can expect to be promising. We explain how the framework shifts our attention in this way, how these new questions differ from past debates about foundations, and sketch one path forward for a philosophy of physics focused on these new questions.
Again, our ultimate goal is to provide genuinely useful and usable guidance on the construction of QFT, however meager or provisional. The construction of QFT is the product of many choices and much reasoning about them. We emphasize that such reasoning is already taking place in support of choices that we actually face as we strive to bring greater specificity to our theoretical knowledge. This reasoning is consequential, whether it takes place in print, project selection, resource allocation, or post-conference debate. It is, therefore, reasoning worthy of attention.